Find four solutions for the equation 3x + 5y = 15.
No book. Thank you very much for your time!!

Found 2 solutions by venugopalramana, funmath:
Answer by venugopalramana(3286)

Find four solutions for the equation 3x + 5y = 15.
No book. Thank you very much for your time!!
GIVE ANY VALUE FOR X AND FIND Y
FOR EX..PUT X=0...SO 3*0+5Y=15...OR..5Y=15..OR...Y=15/5=3
PUT X = 5.....SO 3*5+5Y=15...OR...5Y=0....OR....Y=0
LIKE THIS YOU CAN FIND NY NUMBER OF SOLUTIONS

Answer by funmath(2933)

Find four solutions for the equation 3x + 5y = 15.
:
You get to pick any points you want so make it easy on yourself. Because this is in standard form, (Ax+By=C), the x and y intercepts are easiest to solve for.
:
Let x=0 and solve for y (y-intercept)
3(0)+5y=15
0+5y=15
5y=15
5y/5=15/5
y=3
One solution is: (0,3)
:
Let y=0 and solve for x. (x intercept)
3x+5(0)=15
3x+0=15
3x=15
3x/3=15/3
x=5
Your second solution is (5,0)
:
The next two points, I'm going to pick are letting x=-10 and x=10. I know that they will work out to be whole numbers, it takes practice to find the easiest points.
Let x=-10 and solve for y.
3(-10)+5y=15
-30+5y=15
30-30+5y=15+30
5y=45
5y/5=45/5
y=9
Your third solution is: (-10,9)
:
Let x=10 and solve for y.
3(10)+5y=15
30+5y=15
-30+30+5y=-30+15
5y=-15
5y/5=-15/5
y=-3
Your fourth solution is: (10,-3)
:
Can you think of any others??? (Pick x's that are divisible by 5 to get whole numbers.)
Happy Calculating!!!!